jigglypuff314 (jigglypuff314):

If (w) workers can finish a job in (h) hours, assuming a constant rate, how many hours would it take to complete the job with 1 less worker in terms of (w) and (h)?

jigglypuff314 (jigglypuff314):

Options... if you need it A. \[\frac{ w }{ w-1 }\]B. \[\frac{ (w-1)h }{ w }\]C. \[\frac{ w(w-1) }{ h }\]D.\[\frac{ h }{ w(w-1) }\]E. \[\frac{ wh }{ w-1 }\]

OpenStudy (anonymous):

try doing it with numbers first, rather than \(2\) workers and \(h\) hours i.e. try it with say \(4\) workers take \(5\) hours to do the job, how long would it take \(3\) workers

OpenStudy (raden):

it is an inverse proportional w ---> h w-1 ---> x solve for x : w/(w-1) = x/h

jigglypuff314 (jigglypuff314):

i.e. try it with say 4 workers take 5 hours to do the job, how long would it take 3 workers in the case of ^ would it be 20/3 = 6.66hr?

OpenStudy (anonymous):

\(4\) workers take \(5\) hours to do the job their combines rate is \(\frac{4}{5}\) of a job per hour, so each worker does \(\frac{1}{20}\) of the job per hour therefore \(3\) workers will do \(\frac{3}{20}\) of a job per hour, it will take them \(\frac{20}{3}\) hours to complete the job your answer is right, except i wouldn't use the decimal, leave it as a fraction

jigglypuff314 (jigglypuff314):

ohhh, okay, thank you! :)

OpenStudy (anonymous):

now replace the \(4\) by a \(w\) and the \(5\) by an \(h\) and redo exactly what you did above, using letters instead of numbers in any case that is how i solved it

OpenStudy (phi):

radeng has the right idea. To add some details: It take w people working h hours to finish 1 job in other words, 1 job requires w*h man-hours if you have w-1 people, working for an unknown amount of time t (w-1)*t = 1 job or (w-1) *t = w*h solve for t: t = w h/(w-1)

jigglypuff314 (jigglypuff314):

Thanks you all! :D your comments helped me a lot <3 I wish I could give you all medals :3

OpenStudy (ybarrap):

All explanations above are fantastic. Just one additional note. Look at your options. The units should be "hours.". Also, with 1 less worker, the amout of time to do the job should be "longer." There is only one option that makes sense. This is using a different kind of logic that the ones here, but sometimes helpful for tests.